OpenStudy (anonymous):

bisects DGH so that mDGI is x – 3 and mIGH is 2x – 13. Find the value of x

OpenStudy (anonymous):

Sorry forgot - can you please write out the steps as an explanation ? I don't want just the answer

mathslover (mathslover):

I appreciate your spirit for steps but... I think the question is wrong some how.. Can you draw the diagram?

hartnn (hartnn):

is that GI bisects DGH.......

hartnn (hartnn):

if it is, then those two angles are congruent, x-3 = 2x-13 just solve for x from here....

OpenStudy (anonymous):

Line -> G1 bisects < DGH so that m <DGI is x-3 and m < IGH is 2x - 13 ; find the value of x -Again , sorry , didn't realize that the symbols didn't show up when I copy + pasted =.=

mathslover (mathslover):

http://openstudy.com/study#/updates/505a5048e4b0cc122893a5bf here also it is not written exactly

mathslover (mathslover):

Ok then @hartnn is absolutely correct: \[\large{x-3=\color{blue}{2x-13}}\] \[\large{x-2x=\color{blue}{-13+3}}\] \[\large{-x=\color{blue}{-10}}\] \[\large{x=\color{blue}{10}}\]

hartnn (hartnn):